### v1

# Superconductivity, Anomalous Hall Effect, and Stripe Order in Rhombohedral Hexalayer Graphene

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
Superconductivity, Anomalous Hall Effect, and Stripe Order in Rhombohedral Hexalayer Graphene
Erin Morissette1    Peiyu Qin1    HT Wu1    K. Watanabe2    T. Taniguchi3    J.I.A. Li1 jia_li@brown.edu 1Department of Physics, Brown University, Providence, RI 02912, USA 2Research Center for Functional Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 3International Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(April 7, 2025)
We report the discovery of a unique superconducting phase in rhombohedral hexalayer graphene characterized by its simultaneous emergence with both the anomalous Hall effect and stripe charge order. The onset of stripe charge order is revealed through angle-resolved transport measurements, which show thermally activated insulating behavior along one axis and highly conductive transport along the orthogonal direction. Superconductivity develops exclusively along the high-conductivity axis, giving rise to a one-dimensional–like superconducting channel. This superconducting state exhibits first-order transitions under an out-of-plane magnetic field, consistent with a chiral order parameter that breaks time-reversal symmetry. Most remarkably, thermally driven superconducting transitions display pronounced hysteresis—an uncommon phenomenon that reflects the complex interplay among stripe formation, broken time-reversal symmetry, and superconductivity. Together, these results uncover a previously unidentified quantum phase: a chiral superconductor embedded within an anomalous Hall crystal.
Figure 1: Low-temperature phase diagram of rhombohedral hexalayer graphene. (a) Color-scale map of the longitudinal resistance $R_{\parallel}$, measured at $T=40$ mK and $B=0$, as a function of carrier density $n$ and displacement field $D$. (b) Schematic illustrating the emergent orders derived from panel (a), with superconducting pockets labeled SC i, SC ii, and SC iii. (c) $R$$T$ curves measured from SC i (top), SC ii (middle), and SC iii (bottom). (d) Angle-resolved transport measured at $T=30$ K (top), $20$ K (middle), and $10$ K (bottom). Solid curves denote the best fits using Eqs. (M1) and (M2).
Figure 2: The stripe order and extreme anisotropy in superconducting transport. (a) Resistance versus temperature, $R$$T$, measured with current flowing along different directions relative to the stripe order. The inset shows $R_{\parallel}$, on a linear scale, for current applied parallel (black trace) and slightly misaligned (light blue) from the principal axis. (b) Temperature dependence of the transport anisotropy, defined as $\xi=(R_{\max}-R_{\min})/(R_{\max}+R_{\min})$, which follows a Curie–Weiss form (black dashed line). Top inset: Arrhenius plot of the hard-axis $R$$T$ curve ($\phi=135^{\circ}$), revealing a thermally activated gap $\Delta\approx 4.5\,\text{K}$. Bottom inset: schematic of the stripe charge order, indicating the “easy” (high conductivity) axis at $\phi=225^{\circ}$ and the “hard” (insulating) axis at $\phi=135^{\circ}$. (c), (d) $R$$T$ traces measured in the respective configurations indicated by the insets. When the current flow is partially misaligned with the stripes, the measured resistance is highly sensitive to whether the voltage leads lie along or perpendicular to the stripe direction. As a result, the onset of superconductivity causes negligible voltage drops between leads aligned with the stripes, while leads oriented perpendicular to the stripes still display insulating behavior. All measurements are performed at $n=0.8\times 10^{12}$ cm-2 and $D=958$ mV/nm, inside SC i.
Figure 3: Hysteretic transitions driven by magnetic field and temperature temperature cycling. (a–b) Transport responses as a function of out-of-plane magnetic field $B$: (a) $R_{\perp}$ measured with current flowing parallel to the stripe direction; (b) $R_{\text{cross}}$ measured transverse to the stripes, with current flowing along $\phi=270^{\circ}$. (c–d) Temperature sweep of (c) $R_{\parallel}$ and (d) $R_{\perp}$ measured during both warming and cooling cycles with current applied perpendicular to the stripe orientation at $\phi=135^{\circ}$. (e–f) $R_{\parallel}$ versus $T$ measured during both warming and cooling cycles along (e) $\phi=225^{\circ}$ and (f) $270^{\circ}$, displaying prominent hysteresis in the superconducting transition.
Figure 4: Anisotropic superconducting transport across SC i. (a) Color-scale maps of the longitudinal resistance $R_{\parallel}$ measured with current flowing along the stripes ($\phi=225^{\circ}$, left panel) and transverse to the stripes ($\phi=135^{\circ}$, right panel), plotted as a function of carrier density $n$ and displacement field $D$. Dashed lines outline the boundary of the SC i pocket. White solid lines represent constant-$D$ line cuts analyzed in panels (b–c), while the white dash-dotted line indicates a diagonal cut examined in panel (d). (b–d) Horizontal line cuts at a constant displacement field (b) $D=0.93$, (c) $0.965$, and $0.97$ V/m, showing $R_{\parallel}$ (top) and $R_{\perp}$ (bottom). For clarity, the blue trace in top panels is multiplied by a factor of 5. (d) Longitudinal resistance $R_{\parallel}$ measured along the diagonal dashed line shown in panel (a), traversing both superconducting pockets SC i and SC ii.
Chiral superconductivity has been extensively explored in the unique superfluid A-phase of 3He  Vollhardt and Wölfle (1990), which emerges at ultra-low temperatures around $2$ mK in the strong coupling regime. This remarkable state can be simultaneously viewed as a $p$-wave superconductor and an orbital ferromagnet  Ambegaokar et al. (1974); Walmsley and Golov (2012). To reveal this duality, 3He is confined to a slab geometry, where spatial confinement aligns the orbital magnetic moments perpendicular to the plane, producing a net magnetization at the onset of the superfluid A-phase. By investigating the transport response of moving electron bubble, the magnetic moment associated with the superfluid phase was directly detected via the intrinsic magnus force acting on electron bubbles Ikegami et al. (2013). This distinctive transport signature offers direct demonstration of chiral symmetry breaking in the order parameter of the superfluid A-phase of 3He.
More recently, new graphene-based platforms have emerged as fertile ground for exploring superconductivity intertwined with time-reversal symmetry breaking (TRSB). In particular, tetra- and pentalayer rhombohedral graphene exhibit superconductivity in the high-displacement-field regime that coexists with an anomalous Hall effect Han et al. (2024a); Parra-Martinez et al. (2025); Chen and Schrade (2025); Chou et al. (2024); Yang and Zhang (2024); Christos et al. (2025); Yoon et al. (2025). In these systems, orbital ferromagnetism appears to intertwine with the Cooper pairing instability, giving rise to magnetic-field–driven switching between superconducting states, establishing experimental support for a chiral superconducting state.
Unlike the fluid phase of 3He, where chiral order emerges in relative isolation, the low-temperature phase diagrams of two-dimensional (2D) electron systems—particularly those governed by strong Coulomb interactions—often host a rich interplay of coexisting and competing orders. Notably, recent experiments have demonstrated that the emergence of TRSB in rhombohedral multilayer graphene is intimately tied to spontaneous rotational symmetry breaking Morissette et al. (2025). This connection raises the intriguing possibility that the chiral superconducting state in these systems features broken rotational symmetry—a hypothesis we explore in this work.
To investigate this, we leverage recent advances in angle-resolved transport techniques Wu et al. (2017, 2020); Vafek (2023); Chichinadze et al. (2024); Zhang et al. (2024a); Morissette et al. (2025); Zhang et al. (2025), which provide powerful tools for characterizing how intertwined orders emerge and evolve alongside superconductivity. Applying these methods to rhombohedral hexalayer graphene, we uncover the prominent presence of a stripe charge order that coincides with both superconductivity and the anomalous Hall effect.
Central to our observation is the profound impact of stripe charge order on superconducting transport. Below the stripe formation temperature, charge transport becomes thermally activated and insulating along one principal axis, while remaining highly conductive along the orthogonal direction. Remarkably, at lower temperatures, the superconducting phase inherits this extreme anisotropy: dissipationless current is confined to the high-conductivity axis, forming a one-dimensional–like superconducting channel. By demonstrating the intertwined emergence of superconductivity, the anomalous Hall effect, and stripe charge order, our findings establish a highly unconventional setting for exploring topological superconductivity in low-dimensional confinements  Read and Green (2000); Qi and Zhang (2011); Nadj-Perge et al. (2014); Hua et al. (2024).
The low-temperature phase diagram of rhombohedral hexalayer graphene hosts three distinct superconducting pockets (Fig. 1a), denoted SC i, SC ii, and SC iii in Fig. 1b. Their superconducting transition is demonstrated by vanishing longitudinal resistance with decreasing temperature, as shown by the $R$$T$ curve in Fig. 1c Although all three pockets are situated near orbital ferromagnetic regimes, SC i stands out for its prominent coexistence with the anomalous Hall effect. Consequently, our investigation is focused on the unconventional behavior of SC i, while also examining the angular symmetry in SC ii and SC iii.
We begin by examining the transport properties in the metallic regime above the superconducting transition temperature. The sample geometry and angle-resolved measurement setup, previously discussed in earlier works  Zhang et al. (2024a); Chichinadze et al. (2024); Morissette et al. (2025); Zhang et al. (2025), are also detailed in the Methods section. Fig. 1d presents the angle dependence of longitudinal and transverse resistance, $R_{\parallel}$ and $R_{\perp}$, measured at $n$ and $D$ values within the SC i regime. At $T=30$ K, the transport response appears nearly isotropic, but anisotropy becomes pronounced as the temperature is lowered. According to Fig. 1d, the principal axes of transport anisotropy, defined by the directions of maximum and minimum resistance, are aligned along $\phi=45^{\circ}$ and $135^{\circ}$, respectively.
Figure 2a shows $R_{\parallel}$ as a function of temperature measured along different $\phi$. The maximum and minimum resistance are observed along the principal axes of transport anisotropy, revealing a divergence at low temperature that signals the emergence of extreme transport anisotropy.
To quantify the strength of anisotropy, we define parameter $\xi=(R_{\max}-R_{\min})/(R_{\max}+R_{\min})$, equivalent to $\Delta R/R_{0}$ in Eq. M1-2  Morissette et al. (2025); Zhang et al. (2025); Vafek (2023). The temperature dependence of $\xi$ follows a Curie-Weiss Law (dashed line in Fig. 2b), suggesting a thermodynamic transition driven by nematic instabilities  Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022). Notably, transport along the principal axis of maximum resistivity (minimum conductivity) exhibits activated insulating behavior (inset of Fig. 2b). A linear fit to the Arrhenius plot reveals an energy gap associated with this insulating behavior, $\Delta\approx 4.5$ K. In stark contrast, charge transport along the orthogonal axis remains highly conductive. Such extreme anisotropy, combined with the activated insulating behavior along one principal axis, is consistent with a stripe charge order, with the easy axis near $\phi=45^{\circ}$ (marked by a blue arrow in the inset of Fig. 2b).
Most remarkably, as superconductivity emerges at low temperature, its transport response inherits the spatial anisotropy of the stripe charge order. This extreme anisotropy is manifested in the strong dependence in transition temperature $T_{c}$ on the relative orientation between current flow and the stripe orientation. The highest $T_{c}$ is observed to be around $1$ K when current is applied along the stripe orientation. In contrast, when the current flow is misaligned from the stripe orientation by $45^{\circ}$, as shown in the inset of Fig. 2a, $T_{c}$ is suppressed by almost a factor of 2. The extreme scenario arises when current is applied transverse to the stripe orientation along $\phi=135^{\circ}$. Here, the insulating behavior persists down to the base temperature of the dilution fridge ($T=30$ mK).
Crucially, the extreme transport anisotropy we observe is robust against variations in measurement setup. Beyond the configurations shown in Fig. 2a, Fig. 2c displays transport responses with the current applied along $\phi=90^{\circ}$. In this arrangement, voltage contacts (A, B) that probe the direction perpendicular to the stripes show a low-temperature insulating response, while contacts (A, B) along the stripe direction exhibit diminishing resistance at low temperature. For simplicity, we will refer to this setup as the cross configuration. A similar observation arises under an alternative configuration uses different source–drain and voltage contacts, as illustrated in Fig. 2d: once again, transport measurement reveals an insulating behavior transverse to the stripes and dissipationless parallel to them.
These observations confirm that the coexistence of insulating and superconducting responses originates from the stripe formation, rather than an artifact of specific contact geometry. Additional configurations in Fig. M9 and Fig. M10 further demonstrate that the same set of source-drain-voltage contacts can yield either an insulating or a superconducting response, depending on their orientation relative to the stripe order—ruling out poor electrical contacts as the cause.
A notable and perhaps defining characteristic of the SC i phase is the prevalence of hysteretic behaviors in transport responses. In the following, we systematically examine these hysteretic transitions, providing a unique window into the complex interplay between spontaneous valley polarization, stripe charge order, and superconductivity.
First, we demonstrate that the superconducting phase undergoes hysteretic transitions driven by an out-of-plane magnetic field ($B_{\perp}$), consistent with recent observations in rhombohedral tetra- and pentalayer graphene Han et al. (2024a). When current flows parallel to the stripe direction, Figures 3a–b reveal clear hysteretic switching between two superconducting ground states as $B_{\perp}$ is varied. This behavior is indicative of a first-order transition between degenerate order-parameter configurations, closely analogous to the A-phase of superfluid 3He—a prototypical chiral superconductor Ikegami et al. (2013). Remarkably, similar $B_{\perp}$-driven hysteretic switching is observed even when the current flows perpendicular to the stripe orientation. As shown in the cross configuration of Fig. 3b, hysteresis remains visible despite the transport response being highly resistive. Collectively, these findings establish that the anomalous Hall effect persists throughout both the superconducting and insulating regimes, unveiling a highly exotic scenario in which superconductivity is embedded within an anomalous Hall crystal.
The presence of a crystalline phase necessitates a first-order melting transition, which manifests as hysteresis due to supercooling and superheating during thermal cycling. Consistent with this expectation, temperature-driven hysteresis is clearly observed when probing the insulating response with current applied perpendicular to the stripe direction. As shown in Figs. 3c–d, both $R_{\parallel}$ and $R_{\perp}$ measured along $\phi=135^{\circ}$ exhibit pronounced hysteresis loops during controlled warming and cooling cycles (see Methods for experimental details). Crucially, the shape and magnitude of these loops cannot be attributed to artifacts in temperature regulation. For example, $R_{\perp}$  exhibits a distinct resistance peak during cooling that is completely absent during warming, underscoring its intrinsic origin.
Even more strikingly, hysteretic behavior also emerges across the superconducting transition itself. As illustrated in Figs.3e–f, the superconducting transition temperature $T_{c}$, defined by the vanishing of $R_{\parallel}$, is significantly suppressed on cooling. Notably, the suppressed $T_{c}$ aligns closely with the resistance peak observed in the insulating response during the same cooling cycle (Fig.3d), revealing a compelling connection between the onset of superconductivity and the underlying stripe order.
The observed hysteresis in the superconducting transition is highly unusual. First-order transitions in superconductors typically occur between distinct superconducting phases with different order parameters, as exemplified by the A- and B-phases of superfluid 3He Vollhardt and Wölfle (1990). In contrast, transitions between the superconducting and normal states are generally expected to be continuous and second-order. In our system, however, several plausible mechanisms could give rise to a first-order superconducting transition:
(i) The melting of the stripe charge order is itself a first-order transition. Given the close proximity in temperature between stripe formation and the onset of superconductivity, their coupling could naturally result in a discontinuous superconducting transition with associated hysteresis.
(ii) The superconducting transition temperature ($T_{c}$) is extremely sensitive to the relative orientation between the applied current and the stripe direction. In this scenario, the hysteresis arises from subtle changes in stripe orientation induced either by thermal cycling above $T_{\text{stripe}}$ or by current-direction switching during the measurement sequence.
(iii) There may exist a temperature-driven transition between two distinct superconducting phases, analogous to the A- and B-phases of superfluid 3He. Supercooling and superheating across such a first-order transition would naturally lead to the observed pronounced hysteresis  Schiffer et al. (1992); Schiffer and Osheroff (1995); Tian et al. (2023).
Each of these scenarios highlights the complex and intertwined nature of stripe order and superconductivity in this system, and calls for further investigation into the microscopic origin of the hysteretic superconducting transition.
Taken together, our observations reveal a remarkable and intricate interplay among multiple emergent electronic orders. The magnetic-field–driven hysteretic switching between distinct superconducting ground states strongly supports the presence of a chiral superconducting phase Han et al. (2024a). Simultaneously, the coexistence of stripe charge order and the anomalous Hall effect offers compelling experimental evidence for an anomalous Hall crystal state Tan and Devakul (2024); Dong et al. (2024); Soejima et al. (2024). Moreover, the coexistence of superconductivity with stripe charge order naturally leads to a spatial modulation of the superconducting order parameter, positioning this system as a promising platform for realizing pair density wave states Agterberg et al. (2020).
Previous studies of rhombohedral pentalayer graphene aligned with hexagonal boron nitride (hBN) have also hinted at an anomalous Hall crystal phase Lu et al. (2024, 2025). Our observations of stripe order coexisting with the anomalous Hall effect within a similar region of the $n$$D$ phase diagram suggests that stripe formation may represent a universal ordering tendency in rhombohedral multilayer graphene under strong displacement fields.
Beyond the anomalous Hall crystal interpretation, the coexistence of stripe order and anomalous Hall transport might also be influenced by disorder-induced localization. While conventional transport measurements alone cannot definitively exclude disorder, our sequential measurements, with rotating current flow direction (see Methods), reveal that the stripe orientation can be reconfigured by applied current—a behavior more consistent with an interaction-driven crystalline order than with static disorder. The possibility of such an emergent crystal phase will undoubtedly motivate future experimental investigations.
The stripe order plays a crucial role shaping the properties of superconducting transport. For instance, an outstanding signature of the superconducting phase in rhombohedral tetra- and pentalayer graphene is its remarkable stability against a large out-of-plane magnetic field ($B_{\perp}$)  Han et al. (2024a). In rhombohedral hexalayer graphene, however, this stability exhibits a strong dependence on the orientation of the applied current relative to the stripe direction. When current flows along the stripes, $R_{\parallel}$ remains vanishingly small even at $B_{\perp}=0.5$ T (dark blue trace in Fig.4a), and the corresponding $R$$T$ curves display a sharp superconducting transition, persisting despite the substantial out-of-plane magnetic field.
In contrast, when current is applied along $\phi=270^{\circ}$—misaligned by approximately $45^{\circ}$ from the stripe direction—the application of $B_{\perp}=-0.3$ T leads to a significant increase in $R_{\parallel}$ at low temperatures. Nevertheless, the onset of superconductivity, as indicated by a sharp deviation from the normal-state behavior (highlighted by vertical dashed lines in Fig. 4c), remains largely unaffected by the magnetic field.
The observation of robust superconducting transport at $B=0$ along $\phi=270^{\circ}$ raises an intriguing question about the nature of inter-stripe coupling. As shown in the distinct $R$$T$ behavior in Figs. 4b–c, superconductivity first emerges along the stripe direction near $1$ K and remains robust against $B_{\perp}=-0.3$ T. These intra-stripe channels govern the initial onset of superconductivity. In contrast, the vanishing of $R_{\parallel}$ along $\phi=270^{\circ}$ at low temperatures relies on inter-stripe coupling between adjacent stripes. The application of $B_{\perp}$ primarily disrupts this inter-stripe coupling, leading to a finite $R_{\parallel}$ at low temperatures, while leaving the superconducting onset largely unaffected.
With further increase in $B_{\perp}$, superconductivity is completely suppressed at $B_{\perp}=-1$ T. In this regime, $R$$T$ curves exhibit insulating behavior at low temperatures, irrespective of current flow direction. Notably, the superconducting transition temperature ($T_{c}$) in rhombohedral hexalayer graphene is significantly higher than values reported for tetra- and pentalayer counterparts Han et al. (2024a). This enhanced $T_{c}$ may stem from stronger Coulomb interactions in the flatter bands of hexalayer graphene, as suggested by both theoretical predictions and experimental observations Zhang et al. (2024b); Muten et al. (2021); Pamuk et al. (2017).
The emergence of stripe order is a universal feature across the entire parameter space of the SC i phase. When current flows perpendicular to the stripes, a highly resistive transport response is observed (Fig. 4f), signaling the apparent absence of superconductivity in this configuration. This behavior stands in stark contrast to the vanishing $R_{\parallel}$ observed when current is applied parallel to the stripe orientation (Fig. 4e). The anisotropic nature of the SCi phase is further highlighted by line cuts taken at fixed displacement field $D$ (Fig. 4g; see also Fig. M7).
Even when current is applied transverse to the stripe direction, $R_{\perp}$ drops to zero within the SC i region (Fig. 3h). This observation is consistent with superconductivity developing exclusively along the stripes, which equalizes the potential along their direction, thus resulting in a vanishing $R_{\perp}$ despite a finite, resistive $R_{\parallel}$. Moreover, the extreme anisotropy appears to be unique to the SC i phase. Fig. M8 shows the transport response measured along a diagonal line crossing both the SC i and SC ii phases. While $R_{\parallel}$ exhibits a pronounced discrepancy between the directions parallel and perpendicular to the stripe order in SC i, this discrepancy vanishes in SC ii. The transport anisotropy observed in SC ii is discussed in more detail in the Methods section.
Overall, we employed angle-resolved transport measurements to reveal an unusual superconducting phase embedded in an anomalous Hall crystal. Although this unconventional coexistence has not been explicitly predicted, several theoretical studies have examined Coulomb-driven instabilities involving valley degrees of freedom, where the simultaneous breaking of time-reversal and rotational symmetries promotes spontaneous condensation in momentum space Dong et al. (2023); Jung et al. (2015); Huang et al. (2023); Mandal and Fernandes (2023); Parra-Martinez et al. (2025). Such momentum-space condensation has been widely studied in various multilayer graphene systems Zhang et al. (2024a); Lin et al. (2023), and was recently identified in the multiferroic phase of rhombohedral hexalayer graphene near zero displacement field ($D\approx 0$Morissette et al. (2025).
The emergence of superconductivity from a momentum-space condensed parent state enables Cooper pairing with finite momentum, a phenomenon of great interest due to its direct connection to the superconducting diode effect in the absence of an external magnetic field Lin et al. (2022); Scammell et al. (2021); Yuan and Fu (2021). Given its unique combination of broken symmetries, the SC i phase provides an ideal platform for investigating nonreciprocal superconducting transport at zero magnetic field. While a detailed study of the superconducting diode effect is beyond the scope of this work, it is worth noting that previous reports of a zero-field diode effect also identified the coexistence of superconductivity and a density wave Lin et al. (2022). However, in those cases, the angular symmetry of the density wave could not be resolved due to the constraints imposed by the Hall-bar sample geometry.
Lastly, we highlight the specific correlation between the angular symmetry of superconducting transport and the anisotropy of the metallic phase: the most robust superconducting response consistently occurs along the direction of maximum conductivity (minimum resistivity) observed in the metallic phase above $T_{c}$. While SC ii and SC iii do not exhibit the extreme anisotropy seen in SC i, they share the same angular interplay between superconducting and metallic transport (see Fig. M11). This behavior stands in stark contrast to recent findings in magic-angle twisted trilayer graphene Zhang et al. (2025), where superconductivity is strongest along the direction of minimum conductivity (maximum resistivity) in the normal state. Taken together, these observations establish a fundamental distinction between the superconducting phases in rhombohedral hexalayer graphene and those previously reported in magic-angle graphene moiré systems Cao et al. (2018); Yankowitz et al. (2019); Lu et al. (2019); Park et al. (2021); Hao et al. (2021); Liu et al. (2022).
acknowledgments
J.I.A.L. wishes to acknowledge helpful discussion with Dima Feldman, Oskar Vafek, Aaron Hui, Mathias Scheurer, and Chunli Huang. E.M. and J.I.A.L. acknowledge support from U.S. National Science Foundation under Award DMR-2143384. J.I.A.L. acknowledge partial support from the Air Force Office of Scientific Research. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.