### v2

# Superconductivity, Anomalous Hall Effect, and Stripe Order in Rhombohedral Hexalayer Graphene

thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
Superconductivity, Anomalous Hall Effect, and Stripe Order in Rhombohedral Hexalayer Graphene
Erin Morissette1    Peiyu Qin1    Hai-Tian Wu1    Naiyuan J. Zhang1    K. Watanabe2    T. Taniguchi3    J.I.A. Li1 jia_li@brown.edu 1Department of Physics, Brown University, Providence, RI 02912, USA 2Research Center for Functional Materials, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan 3International Center for Materials Nanoarchitectonics, National Institute for Materials Science, 1-1 Namiki, Tsukuba 305-0044, Japan
(April 16, 2025)
Unconventional superconducting phases are distinguished by broken symmetries in their order parameters. Here, we report the discovery of an exotic superconducting phase in rhombohedral hexalayer graphene at high displacement fields, marked by its coexistence with both a stripe charge order and the anomalous Hall effect. In angle-resolved transport measurements, the onset of stripe order is manifested as thermally activated insulating behavior along one axis, with highly conductive transport along the orthogonal direction. Upon cooling, superconductivity emerges exclusively along the high-conductivity axis, forming one-dimensional–like superconducting channels. This anisotropic superconducting phase exhibits multiple first-order hysteretic transitions. Pronounced thermal hysteresis-between warming and cooling across the superconducting transition-suggests a melting transition of the underlying stripe phase. Additionally, magnetic-field–induced switching—similar to phenomena reported in tetra- and pentalayer graphene—supports the chiral nature of the superconducting state. Together, these findings identify a previously unrecognized quantum phase: a chiral superconductor embedded within a stripy Hall crystal.
Figure 1: Low-temperature phase diagram of rhombohedral hexalayer graphene. (a) Color-scale map of the longitudinal resistance $R_{\parallel}$, measured at $T=40$ mK and $B=0$, as a function of carrier density $n$ and displacement field $D$. (b) Schematic illustrating the emergent orders derived from panel (a), with superconducting pockets labeled SC i, SC ii, and SC iii. (c) $R$$T$ curves measured from SC i (top), SC ii (middle), and SC iii (bottom). (d) Angle-resolved transport measured at $T=30$ K (top), $20$ K (middle), and $10$ K (bottom). Solid curves denote the best fits using Eqs. (M1) and (M2).
Figure 2: The stripe order and extreme anisotropy in superconducting transport. (a) Resistance versus temperature, $R$$T$, measured with current flowing along different directions relative to the stripe order. The inset shows $R_{\parallel}$, on a linear scale, for current applied parallel (black trace) and slightly misaligned (light blue) from the principal axis. (b) Temperature dependence of the transport anisotropy, defined as $\xi=(R_{\max}-R_{\min})/(R_{\max}+R_{\min})$, which follows a Curie–Weiss form (black dashed line). Top inset: Arrhenius plot of the hard-axis $R$$T$ curve ($\phi=135^{\circ}$), revealing a thermally activated gap $\Delta\approx 4.5\,\text{K}$. Bottom inset: schematic of the stripe charge order, indicating the “easy” (high conductivity) axis at $\phi=225^{\circ}$ and the “hard” (insulating) axis at $\phi=135^{\circ}$. (c), (d) $R$$T$ traces measured in the respective configurations indicated by the insets. When the current flow is partially misaligned with the stripes, the measured resistance is highly sensitive to whether the voltage leads lie along or perpendicular to the stripe direction. As a result, the onset of superconductivity causes negligible voltage drops between leads aligned with the stripes, while leads oriented perpendicular to the stripes still display insulating behavior. All measurements are performed at $n=0.8\times 10^{12}$ cm-2 and $D=958$ mV/nm, inside SC i.
Figure 3: Temperature-driven hysteretic transitions and stripe melting. (c–d) Temperature sweep of (c) $R_{\parallel}$ and (d) $R_{\perp}$ measured during both warming and cooling cycles with current applied perpendicular to the stripe orientation at $\phi=135^{\circ}$. (e–f) $R_{\parallel}$ versus $T$ measured during both warming and cooling cycles along (e) $\phi=225^{\circ}$ and (f) $270^{\circ}$, displaying prominent hysteresis in the superconducting transition. Measurements are performed at $D=958$ mV/nm and $n=8\times 10^{12}$ cm-2.
Figure 4: Magnetic-field-driven hysteretic transitions and chiral superconductivity. (a–f) $R_{\parallel}$ (top panels) and $R_{\perp}$ (bottom panels) as a function of out-of-plane magnetic field $B$ measured with current flowing parallel to the stripe direction at (a) $T=0.11$ K, (b) $0.48$ K, (c) $0.81$ K, (d) $1.40$ K, and (e) $1.80$ K. Measurements in (a–f) are performed at $D=930$ mV/nm and $n=8\times 10^{12}$ cm-2. (g-i) $R_{\parallel}$ as a function of $T$ measured at different values of out-of-plane $B$-field and current flow directions.
The superfluid A-phase of 3He, which emerges at ultra-low temperatures around 2 mK in the strong-coupling regime, provides a prototypical example of chiral superconductivity Vollhardt and Wölfle (1990). In this remarkable state, Cooper pairing occurs with non-zero angular momentum, allowing the phase to be simultaneously understood as a $p$-wave superconductor and an orbital ferromagnet Ambegaokar et al. (1974); Walmsley and Golov (2012). This duality is manifested in a net magnetization associated with the angular momentum of the Cooper pairs. When confined to a slab geometry, spatial confinement aligns the orbital magnetic moments perpendicular to the plane, producing two degenerate configurations with magnetic moments pointing either up or down Ikegami et al. (2013). Most remarkably, applying an out-of-plane magnetic field drives a transition between these configurations by reversing the direction of the magnetic moment—a process that manifests in the transport response of moving electron bubbles. This distinctive transport signature provides direct evidence of the chiral nature of the superfluid A-phase of 3He.
Despite decades of effort to identify a chiral superconductor in a solid-state platform, the superfluid A-phase of 3He remained the only known example until recently. A promising candidate has now emerged in the superconducting phases of tetra- and pentalayer rhombohedral graphene Han et al. (2024a); Parra-Martinez et al. (2025); Chen and Schrade (2025); Chou et al. (2024); Yang and Zhang (2024); Christos et al. (2025); Yoon et al. (2025). In these systems, the Cooper pairing instability appears to be intimately linked to orbital ferromagnetism. The resulting superconducting phase exhibits magnetic-field–driven switching between distinct ground states—closely resembling the hallmark behavior of the 3He A-phase Han et al. (2024a).
Unlike the fluid phase of 3He, where chiral order emerges in relative isolation, the low-temperature phase diagrams of two-dimensional (2D) electron systems—particularly those governed by strong Coulomb interactions—often host a rich interplay of coexisting and competing orders Fradkin et al. (2015); Keimer and Moore (2017); Fernandes et al. (2019); Nuckolls and Yazdani (2024); Wang et al. (2023). Among these, multilayer graphene with rhombohedral stacking has emerged as a particularly fertile platform, where the emergence of superconductivity is intertwined with a broad spectrum of electronic orders Zhou et al. (2021a, b, 2022); Han et al. (2024a); Choi et al. (2025). Notably, recent experiments have revealed that the onset of time-reversal symmetry breaking (TRSB) in rhombohedral multilayer graphene is intimately linked to spontaneous rotational symmetry breaking Morissette et al. (2025). This observation raises the compelling possibility that the chiral superconducting state in these systems intrinsically features broken rotational symmetry—a hypothesis we investigate in this work.
To investigate this possibility, we leverage recent advances in angle-resolved transport techniques  Wu et al. (2017, 2020); Vafek (2023); Chichinadze et al. (2024); Zhang et al. (2024a); Morissette et al. (2025); Zhang et al. (2025), which offer powerful tools for probing how intertwined electronic orders emerge and evolve alongside superconductivity. Applying these methods to rhombohedral hexalayer graphene, we uncover a prominent stripe charge order that coexists with both superconductivity and the anomalous Hall effect.
The low-temperature phase diagram of rhombohedral hexalayer graphene hosts three distinct superconducting pockets, as shown in Fig. 1a, labeled SCi, SCii, and SCiii in Fig. 1b. Their superconducting transitions are evident in the $R$$T$ characteristics displayed in Fig. 1c. While all three pockets are proximate to regimes of orbital ferromagnetism, SCi emerges from a valley-polarized metallic phase and uniquely exhibits a pronounced coexistence with stripe charge order (see Fig. M1 and Fig. M2). Accordingly, our study focuses on the unconventional properties of the SCi phase. For completeness, we also comment on the angular transport behavior observed in SCii and SCiii in the Methods section.
We begin by examining the transport properties in the metallic regime above the superconducting transition temperature. The sample geometry and angle-resolved measurement setup have been described in detail in previous studies Zhang et al. (2024a); Chichinadze et al. (2024); Morissette et al. (2025); Zhang et al. (2025). Figure 1d presents the angular dependence of both the longitudinal and transverse resistances, $R_{\parallel}$ and $R_{\perp}$, measured at carrier density $n$ and displacement field $D$ values corresponding to the SCi regime. At $T=30$,K, the transport response is nearly isotropic, but as the temperature decreases, a pronounced anisotropy emerges. As shown in Fig.1d, the principal axes of transport anisotropy—defined by the directions of maximum and minimum resistance—are aligned along $\phi=45^{\circ}$ and $135^{\circ}$, respectively.
Figure 2a shows $R_{\parallel}$ as a function of temperature, measured at various angles $\phi$. The maximum and minimum resistance values, $R_{\max}$ and $R_{\min}$, are observed when current is applied along the principal axes of the transport anisotropy. The increasing divergence between $R_{\max}$ and $R_{\min}$ at low temperatures signals the emergence of extreme transport anisotropy.
We define the anisotropy parameter as $\xi=(R_{\max}-R_{\min})/(R_{\max}+R_{\min})$, which is equivalent to $\Delta R/R_{0}$ in Eq. M1–M2 Morissette et al. (2025); Zhang et al. (2025); Vafek (2023). The dashed line in Fig. 2b represents the best fit of the temperature dependence of $\xi$ to the Curie–Weiss law. The excellent agreement between the experimental data and the Curie–Weiss fit indicates a thermodynamic transition driven by nematic instabilities Chu et al. (2012); Fernandes et al. (2014); Böhmer et al. (2022).
Notably, transport along the principal axis of maximum resistivity (i.e., minimum conductivity) exhibits thermally activated insulating behavior, as shown in the inset of Fig. 2b. A linear fit to the Arrhenius plot yields an energy gap of approximately $\Delta\approx 4.5$ K. In stark contrast, charge transport along the orthogonal axis remains highly conductive. The emergence of activated behavior along only one direction is an unprecedented observation. It points to the formation of an exotic electronic order that goes beyond conventional nematicity. We propose that our findings arise from the formation of a stripe charge order at low temperature  Fogler et al. (1996); Koulakov et al. (1996); Lilly et al. (1999a, b); Du et al. (1999); Pan et al. (1999), with the easy axis aligned near $\phi=45^{\circ}$ (indicated by the blue arrow in the inset of Fig. 2b).
Central to our observation is the profound impact of stripe charge order on superconducting transport. As superconductivity emerges at low temperatures, its transport response inherits the spatial anisotropy imposed by the stripe order. When current is applied along the stripe orientation, the superconducting transition temperature $T_{c}$ reaches a maximum near 1 K. In contrast, when the current is misaligned by $45^{\circ}$—as shown in the inset of Fig. 2a—$T_{c}$ is suppressed by nearly a factor of two. A more dramatic suppression occurs when current is applied transverse to the stripe orientation, along $\phi=135^{\circ}$, where insulating behavior persists down to the base temperature of the dilution refrigerator ($T=30$ mK). This extreme anisotropy is reminiscent of quantum Hall stripe phases Lilly et al. (1999a, b); Du et al. (1999); Pan et al. (1999), though distinguished here by the emergence of superconductivity along the stripe direction.
Crucially, the extreme transport anisotropy we observe is robust against variations in measurement configuration. For example, Fig. 2c shows an alternative setup with current applied along $\phi=90^{\circ}$. In this arrangement, voltage probes positioned perpendicular to the stripe direction (A, B) display a low-temperature insulating response, while probes aligned along the stripes (A, B) exhibit diminishing resistance as temperature decreases. For simplicity, we refer to this setup as the cross configuration. A similar observation is made using different source–drain and voltage contact pairs, as shown in Fig. 2d: once again, the transport response reveals insulating behavior transverse to the stripes and dissipationless behavior along them.
Additional configurations in Fig. M10 and Fig. M11 further demonstrate that whether the response is insulating or superconducting is solely determined by the orientation relative to the stripe order—ruling out poor electrical contacts as the cause. These observations confirm that the coexistence of insulating and superconducting responses originates from the stripe formation, rather than from artifacts related to specific contact geometry.
In the following, we report a rich array of hysteretic behaviors in transport measurements that underscore the remarkable coexistence of superconductivity, stripe charge order, and the anomalous Hall effect.
Most notably, the emergence of stripe order is accompanied by a crystalline phase transition, evidenced by a first-order melting process with pronounced supercooling and superheating. This transition manifests as temperature-driven hysteresis in the insulating response when current is applied perpendicular to the stripe direction. As shown in Figs. 3c–d, both $R_{\parallel}$ and $R_{\perp}$ measured along $\phi=135^{\circ}$ exhibit clear hysteresis loops during controlled warming and cooling (see Methods for details). Importantly, the magnitude and asymmetry of these loops—such as the resistance peak in $R_{\perp}$ observed only during cooling—offer crucial confirmation for their intrinsic origin and rule out experimental artifacts.
The hysteretic $R$$T$ curves also offer insight into the characteristic temperature associated with stripe formation. As shown in Figs. 3a–b, we define the stripe melting temperature ($T_{\text{stripe}}$) as the onset of temperature-driven hysteresis in the insulating response measured perpendicular to the stripe direction. In this context, the resistance cusp observed in $R_{\parallel}$ (Fig. 3a) reflects a thermodynamic transition driven by nematic instability, without involving crystalline order.
Remarkably, $T_{\text{stripe}}$ coincides with the onset of superconductivity when current flows along the stripe direction (Fig. 3c), highlighting a direct link between superconductivity and stripe formation.
Even more striking is the observation of hysteresis across the superconducting transition itself. As shown in Figs. 3c–d, the superconducting transition—defined by the vanishing of $R_{\parallel}$—occurs at a significantly lower temperature during cooling, indicative of a first-order transition.
The presence of hysteresis in the superconducting transition is highly unusual. First-order transitions in superconductors typically occur between distinct superconducting phases with different order parameters, such as the A- and B-phases of superfluid ${}^{3}\mathrm{He}$ Vollhardt and Wölfle (1990). In contrast, transitions between superconducting and normal states are generally expected to be continuous and second-order. In our system, however, several plausible mechanisms may lead to a first-order superconducting transition:
  1. (i)
    The melting of the stripe charge order is itself a first-order transition. Therefore, the hysteresis observed in the superconducting transition may naturally result from the supercooling and superheating associated with stripe melting.
  2. (ii)
    The superconducting transition temperature ($T_{c}$) is extremely sensitive to the relative orientation between the applied current and the stripe direction. In this scenario, the hysteresis arises from subtle changes in stripe orientation induced either by thermal cycling above $T_{\text{stripe}}$ or by current-direction switching during the measurement sequence.
  3. (iii)
    There may exist a temperature-driven transition between two distinct superconducting phases, analogous to the A- and B-phases of superfluid 3He. Supercooling and superheating across such a first-order transition would naturally lead to the observed pronounced hysteresis  Schiffer et al. (1992); Schiffer and Osheroff (1995); Tian et al. (2023).
Each of these scenarios highlights the complex and intertwined nature of stripe order and superconductivity in this system, and calls for further investigation into the microscopic origin of the hysteretic superconducting transition.
Next, we show that the superconducting phase undergoes hysteretic transitions driven by an out-of-plane magnetic field ($B_{\perp}$), indicative of a first-order transition between degenerate order-parameter configurations—closely analogous to the A-phase of superfluid 3He, a prototypical chiral superconductor Ikegami et al. (2013). Figures 4a–c present superconducting transport responses with current flowing parallel to the stripe direction, measured as $B_{\perp}$ is swept back and forth at different temperatures. Both $R_{\parallel}$ and $R_{\perp}$ remain vanishingly small across most of the $B$ range, while field-driven reversals of the underlying valley polarization are marked by prominent peaks in $R_{\perp}$.
Such magnetic-field–driven switching is consistent with recent observations in rhombohedral tetra- and pentalayer graphene Han et al. (2024a), and offers strong support for the chiral nature of the superconducting phase.
While the field-driven reversal of valley polarization manifests as pronounced peaks in $R_{\perp}$, $R_{\parallel}$ remains largely featureless below the stripe formation temperature ($T<0.8$ K). It has been proposed that a resistive signal in the superconducting phase arises when a domain boundary lies between voltage probes Han et al. (2024a). In this context, the contrasting behavior between $R_{\parallel}$ and $R_{\perp}$ may reflect a tendency for domain boundaries to align preferentially along the stripe direction. As temperature increases, field-driven reversal becomes visible in both $R_{\parallel}$ and $R_{\perp}$ around $T=0.8$ K, consistent with the stripe melting temperature identified in Fig. 3. In this regime, domain boundaries are no longer pinned to the stripe orientation.
The magnetic-field–driven hysteresis loop persists up to $T=1.4$ K, indicating that spontaneous valley polarization is a more robust order than both superconductivity and stripe formation. Intriguingly, the onset of the anomalous Hall effect coincides with the nematic transition temperature $T_{\text{nem}}$ (Fig. 3). This connection is further evidenced by the overlapping phase space where extreme transport anisotropy and the anomalous Hall effect coexist, as shown in Fig. M1. A similar correlation has been observed in the multiferroic phase of rhombohedral hexalayer graphene near zero displacement field ($D\approx 0$Morissette et al. (2025). These observations point to a potential common origin for time-reversal and rotational symmetry breaking.
Such a shared origin could naturally arise from Coulomb-driven instabilities of a valley-polarized Fermi surface, which favor spontaneous condensation in momentum space Dong et al. (2023); Jung et al. (2015); Huang et al. (2023); Mandal and Fernandes (2023); Parra-Martinez et al. (2025). Momentum-space condensation has been widely explored in the context of multilayer graphene systems Zhang et al. (2024a); Lin et al. (2023). From a momentum-polarized parent state, the emergence of superconductivity permits finite-momentum Cooper pairing—a scenario directly linked to the superconducting diode effect at zero magnetic field Lin et al. (2022); Scammell et al. (2022); Yuan and Fu (2021). With its distinct combination of broken symmetries, the SC i phase presents an ideal platform to explore nonreciprocal superconducting transport. While a full study of the diode effect lies beyond the scope of this work, we note that prior zero-field diode reports also involved coexisting superconductivity and a density wave Lin et al. (2022). However, in those cases, the angular symmetry of the density wave could not be resolved due to the constraints of Hall-bar geometries.
The existence of valley polarization above the superconducting transition also distinguishes this system from the A-phase of superfluid 3He, where the normal fluid phase preserves time-reversal symmetry.
Taken together, our observations reveal a remarkable and intricate interplay among multiple emergent electronic orders. Above the superconducting transition temperature, the anomalous Hall effect emerges concurrently with pronounced transport anisotropy at the nematic transition $T_{\text{nem}}$, highlighting a metallic phase characterized by spontaneous momentum-space condensation. In the low-temperature regime, magnetic-field–driven hysteretic switching between degenerate superconducting states strongly supports the chiral nature of the superconducting order parameter Han et al. (2024a). The simultaneous onset of superconductivity and stripe charge order naturally leads to a spatial modulation of the superconducting gap, establishing this system as a promising platform for realizing pair density wave states Agterberg et al. (2020). Furthermore, the coexistence of stripe order and the anomalous Hall effect provides compelling experimental evidence for a stripy anomalous Hall crystal Tan and Devakul (2024); Dong et al. (2024a); Soejima et al. (2024); Dong et al. (2024b); Patri et al. (2024); Dong et al. (2024c).
Despite its coexistence and competition with stripe order, the superconducting transition temperature ($T_{c}$) in rhombohedral hexalayer graphene is significantly higher than values reported for other rhombohedral multilayer graphene systems Han et al. (2024a); Zhou et al. (2021b, 2022). This enhanced $T_{c}$ may arise from stronger Coulomb interactions associated with the flatter bands in hexalayer graphene, as suggested by both theoretical predictions and experimental observations  Zhang et al. (2024b); Muten et al. (2021); Pamuk et al. (2017).
Furthermore, previous studies of rhombohedral pentalayer graphene aligned with hexagonal boron nitride (hBN) have also hinted at the presence of an anomalous Hall crystal phase Lu et al. (2024, 2025). Our observation of stripe order coexisting with the anomalous Hall effect within a similar region of the $n$$D$ phase diagram suggests that stripe formation may represent a universal ordering tendency in rhombohedral multilayer graphene under strong displacement fields.
An outstanding signature of the superconducting phase in rhombohedral tetra- and pentalayer graphene is its remarkable stability against large out-of-plane magnetic fields ($B_{\perp}$an et al. (2024a). In rhombohedral hexalayer graphene, however, this stability depends sensitively on the orientation of the applied current relative to the stripe direction. When current flows along the stripe orientation, $R$$T$ curves exhibit a sharp superconducting transition that persists even under a substantial out-of-plane magnetic field of $B_{\perp}=0.3$ T (Fig. 4h). As illustrated in Fig. M12, $R_{\parallel}$ measured along the stripe direction remains nearly zero up to approximately $B_{\perp}=0.75$ T.
In contrast, when current is applied along $\phi=270^{\circ}$—misaligned by approximately $45^{\circ}$ from the stripe direction—the application of $B_{\perp}=-0.3$ T fully suppresses dissipationless superconducting transport at low temperatures (Fig. 4h). Under this field, superconductivity is observed only when the current is perfectly aligned with the stripe orientation. When $B_{\perp}$ is further increased to $-1$ T, superconductivity is completely suppressed, independent of the current direction (Fig. 4i). In this high-field regime, the $R$$T$ curves display an upturn at low temperatures, consistent with the reemergence of a stripe-ordered insulating phase.
Lastly, we comment on the angular interplay between superconductivity and the metallic phase. In SC i, the most robust superconducting response consistently occurs along the direction of maximum conductivity (minimum resistivity) observed in the metallic phase above $T_{c}$. While SC ii and SC iii do not exhibit the extreme anisotropy seen in SC i, they share the same angular interplay between superconducting and metallic transport (see Fig. M13). This behavior stands in stark contrast to recent findings in magic-angle twisted trilayer graphene Zhang et al. (2025), where superconductivity is strongest along the direction of minimum conductivity (maximum resistivity) in the normal state. These observations establish a fundamental distinction between the superconducting phases in rhombohedral hexalayer graphene and those previously reported in magic-angle graphene moiré systems Cao et al. (2018); Yankowitz et al. (2019); Lu et al. (2019); Park et al. (2021); Hao et al. (2021); Liu et al. (2022).
acknowledgments
J.I.A.L. wishes to thank Oskar Vafek for continued support and theoretical guidance in developing the novel scheme of angle-resolved transport measurement. J.I.A.L. also acknowledges helpful discussions with Sankar Das Sarma, Dima Feldman, Aaron Hui, Mathias Scheurer, Chunli Huang, and Boris Shklovskii. E.M. and J.I.A.L. acknowledge support from U.S. National Science Foundation under Award DMR-2143384. J.I.A.L. acknowledge partial support from the Air Force Office of Scientific Research. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) and World Premier International Research Center Initiative (WPI), MEXT, Japan. Part of this work was enabled by the use of pyscan (github.com/sandialabs/pyscan), scientific measurement software made available by the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy.